# GraphBench: Next-generation graph learning benchmarking

### v1

@@ 3.3.2 Combinatorial optimization: Learning to predict optimal solution objectives
Combinatorial optimization (CO) lies at the intersection of optimization, operations research, discrete mathematics, and computer science (Kor+2012). The goal is to find a solution from a discrete set that optimizes a given objective under constraints. Formally, a combinatorial optimization problem $I$ is described by a tuple $(\Omega\ziffer@CommaOri F\ziffer@CommaOri w)$:
@@
CO plays a central role in many practical applications, including vehicle routing, scheduling, and resource allocation (paschos2014applications), underlining the importance of solving such optimization problems effectively in practice. Graphs play a central role in CO due to the inherently discrete structure of many problems and the prevalence of network data. Problems such as max-cut and the independent set problem are naturally expressed on graphs. Among the 21 NP-complete problems identified by karp1972reducibility, ten are graph problems; others, such as set covering, can also be formulated in graph terms. Specifically, variable-constraint relationships often induce bipartite graphs, where variables are connected to the constraints in which they appear. These structural patterns provide strong inductive biases for graph learning.
@@
CO problems are particularly well-suited for learning-based research, i.e., their hardness (karp1972reducibility) motivates MPNNs or GNNs as efficient surrogates for solving large instances. Since many are naturally graph-structured, solving them effectively requires expressive architectures (chen2022milp; chen2024sb). Additionally, large synthetic datasets can be generated for robust benchmarking. Traditional methods can yield optimal solutions for small cases and strong approximations for larger ones, supporting both supervised training and generalization studies. The objective value provides a clear quantitative signal, making CO an ideal testbed for advancing graph learning.
@@
Many recent learning-based optimization work rely on synthetic data (toenshoff2021graph; tonshoff2022one; zhang2023let; karalias2021erdos), either RB graphs (rb_graphs), Erdős-Rényi (ER) graphs (erd6s1960evolution), or Barabási-Albert (BA) graphs (albert2002statistical). Some also utilize pre-existing datasets from earlier graph benchmarks, e.g., (karalias2021erdos) uses TUDatasets (Mor+2020), and li2022rethinking uses Planetoid datasets (yang2016revisiting). Earlier efforts have also produced benchmarks for specific graph CO problems. For graph coloring, examples include hard 3-colorability datasets featuring cliques (mizuno2007toward; mizuno2008constructive), queen graphs (Chv04), Latin square graphs (gomes2002completing), $K$-Insertions graphs (CarOlm02a), and $K$-FullIns graphs (CarOlm02b). The DIMACS benchmark collection (dimacs) provides hundreds of real-world networks and problem instances across various CO domains. Still, their heterogeneity and limited number of instances per category make them less suited for modern machine learning pipelines. More recently, nath2024maxcutbench introduced a max-cut benchmark covering a variety of graph classes. Our motivation differs from these prior works. We aim to design a unified set of graph datasets shared across multiple CO tasks, enabling direct comparison between methods in a consistent experimental setting. While some prior work generates random instances for individual problems, we standardize the process by fixing both the generation procedures and the training-validation splits, ensuring that all methods compare on the same instances.
@@
The supervised setting relies on solver-generated solutions, which can be computationally infeasible for large instances. This motivates the unsupervised setting, which is valuable when ground-truth solutions are not available or expensive to obtain, and directly targets solving the CO problem. We provide a differentiable surrogate loss function $\mathcal{L}\colon\mathbb{R}^{|\Omega(I)|}\to\mathbb{R}$ as well as decoders $d\colon\mathbb{R}^{|\Omega(I)|}\to F(I)$ for each CO problem $I$. The model $m\colon\mathcal{I}\to\mathbb{R}^{|\Omega(I)|}$ is trained in an unsupervised fashion to predict a score for each variable that indicates whether it belongs to the solution set, minimizing
@@
Given a problem defined on a graph $I=G=(V(G)\ziffer@CommaOri E(G))$, the MIS problem is to find the largest subset of non-adjacent nodes, formally,
@@
To construct the datasets, we first define how the input graphs are generated. We consider three well-established random graph models: RB graphs, ER graphs, and BA graphs. These models are widely adopted in machine learning for solving CO problems. The diversity and difficulty of each dataset are controlled via graph generation parameters. Specifically,
@@
We generate 50 000 instances for each dataset and provide both small-scale and large-scale variants. Table 11 shows a summary of the parameters used for graph generation. For RB graphs, parameters are taken from zhang2023let; sanokowski2024diffusion; wenkel2024towards; sun2022annealed. Parameters for large ER graphs are from bother2022what; qiu2022dimes; sun2023difusco; yu2024disco. Parameters for large BA graphs are from bother2022what. The parameters for ER-small and BA-small are adjusted to match the graph size of RB-small. When a parameter is specified as a range, it is sampled uniformly at random for each instance. The only exception to this is the number of nodes in RB graphs, which is determined as described above.
@@
For the supervised setting, we benchmark our baselines on the MIS problem. The task is to predict the CO objective value for the MIS problem. We train all models using the MAE as the loss function and report the final test MAE for evaluation. For the unsupervised setting, we evaluate on all three CO problems, i.e., MIS, max-cut, and graph coloring. The models are trained in an unsupervised fashion using a problem-specific surrogate loss function. At test time, the model’s output scores are converted into a feasible solution to the CO problem using a decoder. The CO problem’s objective function is then used as a metric to measure the quality of the solutions obtained this way. We conduct experiments on small and large synthetic graphs from the RB, ER, and BA families. We compare four baseline models: GIN, GT, MLP, and DeepSets, using the hyperparameters listed in Table 26. Our proposed model uses the encoder-decoder architecture. The encoder, an MLP, processes node features derived from RWSE (Dwivedi2022RWSE) and node degrees. The processor module is the only part that differs between baselines and performs the bulk of the calculations. It is followed by a decoder that generates node representations. For the supervised setting, we apply summation aggregation to obtain a scalar graph-level prediction as the final output. For the unsupervised setting, we leave the node-level scores as the output.
@@ Results
The complete results for the supervised learning task are presented in Table 12. Among the baselines, GIN achieves the best performance across most datasets. We attribute this to MPNNs’ strong inductive bias, which aligns well with the structure of graph-based CO problems. DeepSet generally outperforms the MLP baseline, except for the ER graphs. This suggests that global information aggregation is beneficial for this task. Conversely, the GT performs poorly on some datasets. We hypothesize that this is due to the training difficulties associated with this task. Table 13 shows the baseline results for unsupervised CO. We report the average objective value of solutions obtained by running the decoder on the model’s output scores for test-set problem instances. Out of the four baselines, GIN performs best across all three CO problems, with some exceptions. As expected, the graph-based models GIN and GT generally perform better than MLP and DeepSet. One exception is graph coloring on the (very sparse) BA graphs, where DeepSet outperforms the other two models on both graph sizes. Note that our baselines do not utilize diffusion or other advanced architectures or techniques tailored explicitly for CO. Their performance is therefore much weaker compared to methods that do use them, e.g. sanokowski2024diffusion; zhang2023let. Also note that most learning-based methods currently cannot compete with exact solvers and some hand-crafted heuristics in terms of solution quality, but are often much quicker (sun2023difusco; sanokowski2024diffusion; zhang2023let).
@@ 3.3.3 Algorithmic reasoning: Learning to simulate algorithms
In addition to SAT, many real-world tasks depend on the efficient solution of graph algorithms. Given the active exploration of the intersection of algorithms and neural networks across domains (Esterman2024PUZZLES; Fan2024NPHardEval; Kaiser2015NeuralGPUs; Zaremba2014Learning), our goal is to provide dedicated datasets for graph algorithmic reasoning. Prior work (Velickovic2020NeuralExecution; Xu2020nnreasonabout; Bounsi2024TransfmeetNeuralAlgoReas) and benchmarks such as CLRS (Velickovic2022CLRSbenchmark) have demonstrated that GNN architectures can achieve strong performance on graph problems when provided with hints and graph invariants—tasks that require substantial theoretical expressiveness from graph learning models (Arvind2020WLsubgraph). To broaden this research area, we contribute large-scale datasets for several common graph algorithm tasks, expanding the landscape of neural algorithmic reasoning benchmarks.
@@
We use the encoder-processor-decoder pipeline described in Appendix˜C. For node and graph-level tasks, the pipeline is applied directly with the task-specific encoder described below. Since the graph transformer baseline expects node-level tokenization, we provide an additional graph transform. We leverage the following edge transform which allows us to provide node level tokenization on a modified graph $G^{\prime}$ obtained from the original graph $G$, with $V(G^{\prime})\coloneqq\{(v\ziffer@CommaOri v)\mid v\in V(G)\}\cup E(G)$ and $E(G^{\prime})\coloneqq\{((u\ziffer@CommaOri v)\ziffer@CommaOri(w\ziffer@CommaOri z))\mid u=w\vee u=z\vee v=w\vee v=z\}$.
@@ Results
We report results on all datasets for each of the three selected difficulties. These are denoted using the terms easy, medium, and hard to reflect the additional generalization requirements introduced by out-of-distribution sampling. For classification tasks, we report F1 scores, and for regression tasks, we report MAE. All results are grouped by their respective estimated task difficulty in Table˜16 and Table˜17 below. Additionally, we report size generalization results using pre-trained baselines from the medium setting of our training procedure. For this, we evaluate each pre-trained model from 128 to 512 graph node counts. Results are provided in Table˜18. For minimum spanning tree, max clique and max matching datasets show improved performance with the graph transformer baseline. However, on the max flow, bridges, and Steiner tree datasets, the GIN baseline model performs better across the proposed difficulty levels. Furthermore, we observe a reduction in F1 scores with increasing difficulty across datasets, highlighting the varying generator selections in the training data. However, the differences are not as pronounced as expected for the bridges dataset. We further observe that the GIN baseline is inconsistent across seeds for the minimum spanning tree and Steiner tree tasks, whereas the graph transformer baseline does not exhibit similar issues. Overall, the performance of both baselines is the lowest for Steiner trees and max clique tasks, indicating that these tasks are more complex for graph learning baselines to learn. Similar to the results obtained in the CLRS benchmark (Velickovic2022CLRSbenchmark) and by 2024muellertowards for an expressive graph transformer architecture, MST, bridges, and topological sorting provide more manageable tasks for graph learning baselines. We note that, despite previous benchmarks such as CLRS, there are currently no other baseline results on the synthetic algorithmic reasoning datasets we provide.
@@
Table 18: Size generalization results on algorithmic reasoning datasets from Section˜3.3.3. Each column represents an evaluation on 1000 graphs with the given number of nodes (128–512). For each experiment, the same generation parameters were used as in the results presented in Table˜16 and Table˜17. OOT denotes the case where computation of the underlying size generalization data was not completed in less than 24 hours of compute on a single cluster node. All generated graphs are based on the medium difficulty setting. We use the same single seed for each pretrained model from each task to provide inference results on different sizes.
@@ 3.4.1 Weather forecasting: Medium-range atmospheric state prediction
Weather forecasting has long been critical for agriculture, energy, and public safety (diehl2013visual; ramar2014semantic). Traditional numerical weather prediction (NWP) is accurate but relies on computationally expensive physics-based simulations (bauer2015quiet). Recent machine learning methods offer faster, scalable alternatives that can outperform NWP for medium-range forecasts and extreme events—while improving uncertainty quantification (price2023gencast). Earth system data are sparse, irregular, and interconnected (reichstein2019deep), with interactions spanning scales from local convection to planetary waves (bauer2015quiet). Graph-based models naturally capture these multi-scale, non-Euclidean dependencies, enabling flexible representations of local and global atmospheric interactions.
@@
The idea of utilizing machine learning models for weather forecasting first emerged in 2020, when rasp2020weatherbench and weyn2020improving employed CNN architectures to predict global weather at resolutions of 5.625ř and 1.9ř, respectively. A key breakthrough occurred when keisler2022forecasting first applied MPNNs to represent the globe more naturally, enabling the prediction of 3D states 6 hours ahead and facilitating multi-day forecasts. This approach has already achieved a skill comparable to that of 1ř global NWP (GFS/ECMWF) on specific metrics, outperforming prior data-driven models. Building on this idea, lam2023learning introduced GraphCast, a multi-scale MPNN-based system that delivers today’s forecasts at 0.25ř resolution. GraphCast achieved unprecedented accuracy, outperforming even ECMWF’s high-resolution deterministic model (HRES) out to ten days. The success of these deterministic forecasts raised the question of probabilistic prediction. price2023gencast introduced GenCast, a generative weather model that produces an ensemble of forecasts rather than a single deterministic run. For this, they used a diffusion-based approach built on an MPNN backbone, outperforming ECMWF’s ensemble system (ENS), the world’s top operational probabilistic forecast on over 97% of evaluated metrics. Several other advances can be noted. oskarsson2024probabilistic have proposed Graph-EFM, a model featuring a hierarchical GNN that handles global and regional forecasts seamlessly. With OneForecast, gao2025oneforecast have introduced an MPNN framework using nested multi-scale graphs and adaptive message passing to improve local extreme event predictions within a global context. These works indicate a trend toward hybrid solutions that densify graphs in target regions for high-resolution detail while maintaining global consistency. Current state-of-the-art models, such as Pangu-Weather (bi2022pangu), Aurora (bodnar2024aurora), and FGN (alet2025skillful), all present excellent results that, in many areas, far exceed those of conventional NWP methods. It is important to note, however, that while machine learning-based approaches exceed NWP skill on tested time scales, they do not explicitly enforce conservation laws. Over long integration horizons, this can lead to subtle physical drifts that may limit specific climate-related uses. This becomes particularly relevant under regime shifts, for example, caused by climate change, that lie outside the distributions the models were trained on. Furthermore, compared to NWP models, ML-based approaches offer very limited interpretability, functioning mainly as black boxes.
@@
We utilize reanalysis data from the ERA5 dataset, which has been preprocessed via the WeatherBench2 pipeline (rasp2024weatherbench). Multiple resolutions of this dataset are available. We use a downsampled version containing a $64\times 32$ equiangular grid, employing a conservative area-preserving interpolation. ERA5 data has a temporal resolution of six hours, with timesteps at 0h, 6h, 12h, and 18h. Each weather state contains 62 physical and derived variables: 15 are defined across the 13 pressure levels, and 47 are defined at the surface. Of those, we only use five variables for the surface and six for the atmospheric levels each, as done by price2023gencast. The pressure-level variables include quantities such as temperature, humidity, and wind speed. The surface-level variables also include more general information about the location, such as a land-sea mask, sea level pressure, and precipitation.

### v2

@@ 3.3.2 Combinatorial optimization: Learning to predict optimal solution objectives
Combinatorial optimization (CO) lies at the intersection of optimization, operations research, discrete mathematics, and computer science (Kor+2012). The goal is to find a solution from a discrete set that optimizes a given objective under constraints. Formally, a combinatorial optimization problem instance $I$ is described by a tuple $(\Omega\ziffer@CommaOri F\ziffer@CommaOri w)$:
@@
CO plays a central role in many practical applications, including vehicle routing, scheduling, and resource allocation (paschos2014applications), underscoring the importance of solving such optimization problems effectively. Graphs play a central role in CO due to the inherently discrete structure of many problems and the prevalence of network data. Problems such as max-cut and the independent set problem are naturally expressed on graphs. Among the 21 NP-complete problems identified by karp1972reducibility, ten are graph problems; others, such as set covering, can also be formulated in graph terms. Specifically, variable-constraint relationships often induce bipartite graphs, where variables are connected to the constraints in which they appear. These structural patterns provide strong inductive biases for graph learning.
@@
CO problems are particularly well-suited for learning-based research, i.e., their hardness (karp1972reducibility) motivates the use of MPNNs or GNNs as efficient surrogates for solving large instances. Since many are naturally graph-structured, solving them effectively requires expressive architectures (chen2022milp; chen2024sb). Additionally, large synthetic datasets can be generated for robust benchmarking. Traditional methods can yield optimal solutions for small cases and strong approximations for larger ones, supporting both supervised training and generalization studies. The objective value provides a clear quantitative signal, making CO an ideal testbed for advancing graph learning.
@@
Many recent learning-based optimization work rely on synthetic data (toenshoff2021graph; tonshoff2022one; zhang2023let; karalias2021erdos), either RB graphs (rb_graphs), Erdős-Rényi (ER) graphs (erd6s1960evolution), or Barabási-Albert (BA) graphs (albert2002statistical). Some also utilize pre-existing datasets from earlier graph benchmarks, e.g., (karalias2021erdos) uses TUDatasets (Mor+2020), and li2022rethinking uses Planetoid datasets (yang2016revisiting). Earlier efforts have also produced benchmarks for specific graph CO problems. For graph coloring, examples include hard 3-colorability datasets featuring cliques (mizuno2007toward; mizuno2008constructive), queen graphs (Chv04), Latin square graphs (gomes2002completing), $K$-Insertions graphs (CarOlm02a), and $K$-FullIns graphs (CarOlm02b). The DIMACS benchmark collection (dimacs) provides hundreds of real-world networks and problem instances across various CO domains. Still, their heterogeneity and limited number of instances per category make them less suited for modern machine learning pipelines. More recently, nath2024maxcutbench introduced a max-cut benchmark covering a variety of graph classes. Our motivation differs from these prior works. We aim to design a unified set of graph datasets shared across multiple CO tasks, enabling direct comparison between methods in a consistent experimental setting. While some prior work generates random instances for individual problems, we standardize the process by fixing both the generation procedures and the training-validation splits, ensuring that all methods compare on the same instances. Figure˜6 provides an overview of our graph generators, the tasks we perform, and the parameters we chose for generation.
@@
The supervised setting relies on solver-generated solutions, which can be computationally infeasible for large instances. This motivates the unsupervised setting, which is valuable when ground-truth solutions are not available or expensive to obtain, and directly targets solving the CO problem. We provide a differentiable surrogate loss function $\mathcal{L}\colon\mathbb{R}^{|\Omega(I)|}\to\mathbb{R}$ as well as decoders $d\colon\mathbb{R}^{|\Omega(I)|}\to F(I)$ for each CO problem. The model $m\colon\mathcal{I}\to\mathbb{R}^{|\Omega(I)|}$ is trained in an unsupervised fashion to predict a score for each variable that indicates whether it belongs to the solution set, minimizing
@@
Given a problem defined on a graph $G=(V(G)\ziffer@CommaOri E(G))$, the MIS problem is to find the largest subset of non-adjacent nodes, formally,
@@
To construct the datasets, we first define how the input graphs are generated. We consider three well-established random graph models: RB graphs, ER graphs, and BA graphs. These models are widely used in machine learning to solve CO problems. The diversity and difficulty of each dataset are controlled via graph generation parameters. Specifically,
@@
We generate 50 000 instances for each dataset and provide both small-scale and large-scale variants. Table˜11 shows a summary of the parameters used for graph generation. For RB graphs, parameters are taken from zhang2023let; sanokowski2024diffusion; wenkel2024towards; sun2022annealed. Parameters for large ER graphs are from bother2022what; qiu2022dimes; sun2023difusco; yu2024disco. Parameters for large BA graphs are from bother2022what. The parameters for ER-small and BA-small are adjusted to match the graph size of RB-small. When a parameter is specified as a range, it is sampled uniformly at random for each instance. The only exception to this is the number of nodes in RB graphs, which is determined as described above.
@@
For the supervised setting, we benchmark our baselines on the MIS problem. The task is to predict the CO objective value for the MIS problem. We train all models using the MAE as the loss function and report the final test MAE for evaluation. For the unsupervised setting, we evaluate on all three CO problems, i.e., MIS, max-cut, and graph coloring. The models are trained unsupervised using a problem-specific surrogate loss function. At test time, the model’s output scores are converted into a feasible solution to the CO problem using a decoder. The CO problem’s objective function is then used as a metric to measure the quality of the solutions obtained this way. We conduct experiments on small and large synthetic graphs from the RB, ER, and BA families. We compare four baseline models: GIN, GT, MLP, and DeepSets, using the hyperparameters listed in Table˜26. Our proposed model uses the encoder-decoder architecture. The encoder, an MLP, processes node features derived from RWSE (Dwivedi2022RWSE) and node degrees. The processor module is the only part that differs between baselines and performs the bulk of the calculations. It is followed by a decoder that generates node representations. For the supervised setting, we apply summation aggregation to obtain a scalar graph-level prediction as the final output. For the unsupervised setting, we leave the node-level scores as the output.
@@ Results
The complete results for the supervised learning task are presented in Table˜12. Among the baselines, GIN achieves the best performance across most datasets. We attribute this to MPNNs’ strong inductive bias, which aligns well with the structure of graph-based CO problems. DeepSet generally outperforms the MLP baseline, except for the ER graphs. This suggests that global information aggregation is beneficial for this task. Conversely, the GT performs poorly on some datasets. We hypothesize that this is due to the training difficulties associated with this task. Table˜13 shows the baseline results for unsupervised CO. We report the average objective value of solutions obtained by running the decoder on the model’s output scores for test-set problem instances. Out of the four baselines, GIN performs best across all three CO problems, with some exceptions. As expected, the graph-based models GIN and GT generally perform better than MLP and DeepSet. One exception is graph coloring on the (very sparse) BA graphs, where DeepSet outperforms the other two models on both graph sizes. Note that our baselines do not utilize diffusion or other advanced architectures or techniques tailored explicitly for CO. Their performance is therefore much weaker compared to methods that do use them, e.g. sanokowski2024diffusion; zhang2023let. Also note that most learning-based methods currently cannot compete with exact solvers and some hand-crafted heuristics in terms of solution quality, but are often much quicker (sun2023difusco; sanokowski2024diffusion; zhang2023let).
@@
Figure 7: Example graphs and solutions across all graph generators and all performed tasks for the Algorithmic Reasoning datasets with our chosen parameters.
@@ 3.3.3 Algorithmic reasoning: Learning to simulate algorithms
In addition to SAT, many real-world tasks depend on the efficient solution of graph algorithms. Given the active exploration of the intersection of algorithms and neural networks across domains (Esterman2024PUZZLES; Fan2024NPHardEval; Kaiser2015NeuralGPUs; Zaremba2014Learning), our goal is to provide dedicated datasets for graph algorithmic reasoning. Prior work (Velickovic2020NeuralExecution; Xu2020nnreasonabout; Bounsi2024TransfmeetNeuralAlgoReas) and benchmarks such as CLRS (Velickovic2022CLRSbenchmark) have demonstrated that GNN architectures can achieve strong performance on graph problems when provided with hints and graph invariants—tasks that require substantial theoretical expressiveness from graph learning models (Arvind2020WLsubgraph). To broaden this research area, we contribute large-scale datasets for several common graph algorithm tasks, expanding the landscape of neural algorithmic reasoning benchmarks. Figure˜7 provides an overview of our graph generators, the tasks we perform and the parameters we chose for generation.
@@
We use the encoder-processor-decoder pipeline described in Appendix˜B. For node and graph-level tasks, the pipeline is applied directly with the task-specific encoder described below. Since the graph transformer baseline expects node-level tokenization, we provide an additional graph transform. We leverage the following edge transform which allows us to provide node level tokenization on a modified graph $G^{\prime}$ obtained from the original graph $G$, with $V(G^{\prime})\coloneqq\{(v\ziffer@CommaOri v)\mid v\in V(G)\}\cup E(G)$ and $E(G^{\prime})\coloneqq\{((u\ziffer@CommaOri v)\ziffer@CommaOri(w\ziffer@CommaOri z))\mid u=w\vee u=z\vee v=w\vee v=z\}$.
@@ Results
We report results on all datasets for each of the three selected difficulties. These are denoted using the terms easy, medium, and hard to reflect the additional generalization requirements introduced by out-of-distribution sampling. For classification tasks, we report F1 scores, and for regression tasks, we report MAE. All results are grouped by their respective estimated task difficulty in Table˜16 and Table˜17 below. Additionally, we report size generalization results using pre-trained baselines from the medium setting of our training procedure. For this, we evaluate each pre-trained model across 128 to 512 graph nodes. Results are provided in Table˜18. For minimum spanning tree, max clique and max matching datasets show improved performance with the graph transformer baseline. However, on the max flow, bridges, and Steiner tree datasets, the GIN baseline model performs better across the proposed difficulty levels. Furthermore, we observe a reduction in F1 scores with increasing difficulty across datasets, highlighting the varying generator selections in the training data. However, the differences are not as pronounced as expected for the bridges dataset. We further observe that the GIN baseline is inconsistent across seeds for the minimum spanning tree and Steiner tree tasks, whereas the graph transformer baseline does not exhibit similar issues. Overall, the performance of both baselines is the lowest for Steiner trees and max clique tasks, indicating that these tasks are more complex for graph learning baselines to learn. Similar to the results obtained in the CLRS benchmark (Velickovic2022CLRSbenchmark) and by 2024muellertowards for an expressive graph transformer architecture, MST, bridges, and topological sorting provide more manageable tasks for graph learning baselines. We note that, despite previous benchmarks such as CLRS, there are currently no other baseline results on the synthetic algorithmic reasoning datasets we provide.
@@
Table 18: Size generalization results on algorithmic reasoning datasets from Section˜3.3.3. Each column represents an evaluation on 1000 graphs with the given number of nodes (128–512). For each experiment, the same generation parameters were used as in the results presented in Table˜16 and Table˜17. OOT denotes the case where the underlying size generalization data was not computed within 24 hours of computation on a single cluster node. All generated graphs are based on the medium difficulty setting. We use the same single seed for each pretrained model from each task to provide inference results on different sizes.
@@ 3.4 Earth systems
Figure 8: Down-scaled version of our node icosahedron spanning the globe. Each node contains weather variables of the corresponding location across 13 pressure levels.
@@ 3.4.1 Weather forecasting: Medium-range atmospheric state prediction
Weather forecasting has long been critical for agriculture, energy, and public safety (diehl2013visual; ramar2014semantic). Traditional numerical weather prediction (NWP) is accurate but relies on computationally expensive physics-based simulations (bauer2015quiet). Recent machine learning methods offer faster, scalable alternatives that can outperform NWP for medium-range forecasts and extreme events—while improving uncertainty quantification (price2023gencast). Earth system data are sparse, irregular, and interconnected (reichstein2019deep), with interactions spanning scales from local convection to planetary waves (bauer2015quiet). Graph-based models naturally capture these multi-scale, non-Euclidean dependencies, enabling flexible representations of local and global atmospheric interactions. Figure˜8 provides an intuitive image of how the graph icosahedron spans the globe.
@@
The idea of utilizing machine learning models for weather forecasting first emerged in 2020, when rasp2020weatherbench and weyn2020improving employed CNN architectures to predict global weather at resolutions of 5.625ř and 1.9ř, respectively. A key breakthrough occurred when keisler2022forecasting first applied MPNNs to more naturally represent the globe, enabling the prediction of 3D states 6 hours ahead and facilitating multi-day forecasts. This approach has already achieved a skill comparable to that of 1ř global NWP (GFS/ECMWF) on specific metrics, outperforming prior data-driven models. Building on this idea, lam2023learning introduced GraphCast, a multi-scale MPNN-based system that delivers today’s forecasts at 0.25ř resolution. GraphCast achieved unprecedented accuracy, outperforming even ECMWF’s high-resolution deterministic model (HRES) out to ten days. The success of these deterministic forecasts raised the question of probabilistic prediction. price2023gencast introduced GenCast, a generative weather model that produces an ensemble of forecasts rather than a single deterministic run. For this, they used a diffusion-based approach built on an MPNN backbone, outperforming ECMWF’s ensemble system (ENS), the world’s top operational probabilistic forecast on over 97% of evaluated metrics. Several other advances can be noted. oskarsson2024probabilistic have proposed Graph-EFM, a model featuring a hierarchical GNN that seamlessly handles global and regional forecasts. With OneForecast, gao2025oneforecast have introduced an MPNN framework using nested multi-scale graphs and adaptive message passing to improve local extreme event predictions within a global context. These works indicate a trend toward hybrid solutions that densify graphs in target regions for high-resolution detail while maintaining global consistency. Current state-of-the-art models, such as Pangu-Weather (bi2022pangu), Aurora (bodnar2024aurora), and FGN (alet2025skillful), all present excellent results that, in many areas, far exceed those of conventional NWP methods. It is important to note, however, that while machine learning-based approaches exceed NWP skill on tested time scales, they do not explicitly enforce conservation laws. Over long integration horizons, this can lead to subtle physical drifts that may limit specific climate-related uses. This becomes particularly relevant under regime shifts, for example, caused by climate change, that lie outside the distributions the models were trained on. Furthermore, compared to NWP models, ML-based approaches offer very limited interpretability, functioning mainly as black boxes.
@@
We utilize reanalysis data from the ERA5 dataset, which has been preprocessed via the WeatherBench2 pipeline (rasp2024weatherbench). Multiple resolutions of this dataset are available. We use a downsampled version containing a $64\times 32$ equiangular grid, employing a conservative area-preserving interpolation. ERA5 data has a temporal resolution of six hours, with timesteps at 0h, 6h, 12h, and 18h. Each weather state contains 62 physical and derived variables: 15 are defined across the 13 pressure levels, and 47 are defined at the surface. Of those, we only use five variables for the surface and six for the atmospheric levels each, as done by price2023gencast. The pressure-level variables include temperature, humidity, and wind speed. The surface-level variables also include more general information about the location, such as a land-sea mask, sea level pressure, and precipitation.